# Nonlinearly Preconditioned Gradient Methods under Generalized Smoothness

### v1

@@ Abstract
We analyze nonlinearly preconditioned gradient methods for solving smooth minimization problems. We introduce a generalized smoothness property, based on the notion of abstract convexity, that is broader than Lipschitz smoothness and provide sufficient first- and second-order conditions. Notably, our framework encapsulates algorithms associated with the clipping gradient method and brings out novel insights for the class of $(L_{0},L_{1})$-smooth functions that has received widespread interest recently, thus allowing us to go beyond already established methods. We investigate the convergence of the proposed method in both the convex and nonconvex setting.
Keywords:
nonconvex optimization $\cdot$ generalized smoothness $\cdot$ first-order methods
AMS Subject Classification:
65K05 $\cdot$ 49J52 $\cdot$ 90C30
@@ 1 Introduction and preliminaries
where $f$ is a continuously differentiable and possibly nonconvex function. While gradient descent is a reliable solver for this type of problems, in many cases it does not fully take advantage of the cost function’s properties and requires well-tuned or costly stepsize strategies to converge.
In this paper we thus focus on non-linearly preconditioned gradient methods that are tailored to the properties of the cost function. Given stepsizes $\gamma>0$ and $\lambda>0$ and a starting point $x^{0}\in\mathbb{R}^{n}$, we consider the following iteration:
@@
where $\phi^{*}:\mathbb{R}^{n}\to\mathbb{R}$ is convex and is called the dual reference function. The convex conjugate of $\phi^{*}$, $\phi$ is called the reference function and its properties are crucial to our analysis. This general framework was originally analyzed in [maddison2021dual] under the so-called dual relative smoothness condition for convex and essentially smooth functions. In the general nonconvex and composite nonsmooth setting, it was studied in [laude2022anisotropic] under a condition called the anisotropic descent property, which was itself first introduced in [laude2023dualities].
Our analysis is mainly focused on two types of reference functions that are generated by a kernel function $h:\mathbb{R}\to\mathbb{R}_{+}\cup\{\infty\}$, thus resulting in two different families of algorithms. The first one is given by the composition with the Euclidean norm, i.e. $\phi=h\circ\|\cdot\|$ and is referred to as the isotropic reference function. In this case, the main iteration (2) takes the form of gradient descent with a scalar stepsize that depends on the iterates. The second is a separable sum obtained via $\phi(x)=\sum_{i=1}^{n}h(x_{i})$, henceforth called the separable reference function. In this case, (2) becomes gradient descent with a coordinate-wise stepsize that depends on the iterates of the algorithm. We remark that although the separable reference function makes the analysis of the method more convoluted, it generates more interesting algorithms, akin to the ones that are often used in practice.
@@ 1.1 Motivation
Unifying framework for clipping algorithms. Gradient clipping and signed gradient methods have garnered attention in recent years due to their efficiency in neural network training and other applications [bernstein2018signsgd, gorbunov2020stochastic, zhang2019gradient, zhang2020improved, zhang2020adaptive, koloskova2023revisiting, kunstner2024heavy]. The intuition behind gradient clipping is straightforward, since by clipping one does not allow the potentially very large stochastic gradients to hinder the training. Nevertheless, in many cases the clipping threshold and stepsize should be carefully tuned in practice, otherwise leading to suboptimal performance [koloskova2023revisiting]. While algorithms of this type have been analyzed under various smoothness and stochasticity assumptions, there does not seem to exist a simple unifying framework that encapsulates them. Motivated by this gap, we propose a framework that provides further insights into existing methods but also naturally generates new algorithms.
@@
for $\phi=\tfrac{1}{2}\|\cdot\|^{2}$ and $\bar{y}=T_{{L^{-1}},{1}}(\bar{x})=\bar{x}-\tfrac{1}{L}\nabla\phi^{*}(\nabla f(\bar{x}))$. Going beyond the standard Lipschitzian assumptions, it is natural to consider reference functions $\phi$ that generate less restrictive descent inequalities thus allowing us to efficiently tackle more general problems. This is exactly the anisotropic descent property [laude2022anisotropic, Definition 3.1] and minimizing this upper bound leads to the algorithm described in (2) (for $\lambda$ fixed). Considering (3), it is natural to examine reference functions that are strongly convex and grow faster than a quadratic in order to obtain a less restrictive descent inequality. It turns out that in many interesting cases, the preconditioner in (2) becomes similar to a sigmoid function and the algorithmic step takes the form of popular algorithms. We present examples of kernel functions and the corresponding preconditioners in Table 1.
The abstraction discussed in the previous paragraph is in fact tightly connected to the notion of $\Phi$-convexity (also known as $c$-concavity in the optimal transport literature), which states that a function is $\Phi$-convex if it can be written as the pointwise supremum over a family of nonlinear functions. Similarly to the fact that every proper, lsc and convex function can be expressed as a pointwise supremum of its affine minorizers [RoWe98, Theorem 8.13], anisotropic smoothness then requires that $-f$ is a pointwise supremum of nonlinear minorizers. This fact is once again in parallel to classical $L$-smoothness, which requires that $-f$ is the pointwise supremum over concave quadratics, and leads to an envelope representation of $f$ that is useful in studying the corresponding calculus. In that regard, anisotropic smoothness is a straightforward extension of Lipschitz smoothness. A visualization of the concept of $\Phi$-convexity is shown in Figure 1.
@@ 1.2 Our contribution
  • We describe a common nonlinear gradient preconditioning scheme for the main iterates of popular algorithms including gradient clipping, Adam, Adagrad and recently introduced methods for $(L_{0},L_{1})$-smoothness. These preconditioners are gradients of smooth convex functions and have sigmoid shape, reminiscent of common activation functions in neural networks.
  • We introduce $(L,\bar{L})$-anisotropic smoothness, which extends [laude2022anisotropic] allowing for two constants and a reference function $\phi$ of possibly non-full domain and prove that it is less restrictive than $L$-smoothness. Through a novel technique we study necessary and sufficient first- and second-order conditions for $(L,\bar{L})$-anisotropic smoothness, in the process obtaining novel characterizations for the forward operator $T_{{\gamma},{\lambda}}$, which in turn leads to new insights into the class of $(L_{0},L_{1})$-smooth functions.
  • We analyze the convergence of (2) in the nonconvex setting, obtaining results for our stationarity measure. In the convex setting, we prove the sublinear rate of the method for a large family of isotropic reference functions, utilizing only a simple dual characterization. In the more challenging case where $\phi$ is separable, we present an unconventional proof that is based on the envelope representation of anisotropic smoothness. We are thus able to obtain standard $O(1/K)$ convergence rates for large classes of functions.
@@ 1.3 Related work
Dual space preconditioning and anisotropic smoothness. The scheme described in (2) was originally introduced in [maddison2021dual] in the convex setting, where sufficient second-order conditions and the convergence of the method were established. In [laude2022anisotropic] the anisotropic smoothness condition was studied, which was shown to naturally lead to the convergence of the method in the nonconvex and proximal case. Moreover, in [leger2023gradient] the scheme was also analyzed under the general framework of $\Phi$-convexity and a sufficient second-order condition for anisotropic smoothness was provided. Nevertheless, this requires that $\phi\in\mathcal{C}^{4}(\mathbb{R}^{n})$ satisfies the non-negative cross curvature condition from optimal transport (NNCC) (see [figalli2011multidimensional, Assumption (B3)] and [leger2023gradient, Definition 2.8]), which is a strong assumption that does not hold for many interesting reference functions. Recently, the method was also extended to measure spaces in [bonet2024mirror]. Furthermore, a relaxed proximal point algorithm with nonlinear preconditioning akin to (2) for solving inclusion problems was studied in [laude2023anisotropic].
@@
Generalized smoothness. Our work is also connected to other notions of generalized smoothness, i.e. descent inequalities beyond the standard Lipschitzian assumptions. To begin with, Bregman (relative) smoothness is a popular extension of Lipschitz smoothness [bauschke2017descent, lu2018relatively, bolte2018first, ahookhosh2021bregman] that can encapsulate a wide variety of functions such as those whose Hessians exhibit a certain polynomial growth [lu2018relatively]. Other notions of generalized smoothness include Hölder smoothness [bredies2008forward, nesterov2015universal] and also higher-order smoothness where higher-order derivatives of $f$ are Lipschitz continuous [nesterov2006cubic, doikov2020inexact].
@@
Recently, a new concept of smoothness has been introduced in order to capture the cases where the norm of the Hessian is upper bounded by some function of the norm of the gradient [zhang2019gradient, chen2020understanding, li2024convex]. This condition, which in its most popular form is called $(L_{0},L_{1})$-smoothness [zhang2019gradient, Definition 1], has received widespread attention. Various existing methods have been analyzed under this new smoothness condition [wang2023convergence, faw2023beyond, koloskova2023revisiting], while also new ones have been proposed [gorbunov2024methods, vankov2024optimizing]. We remark that a number of these algorithms can actually be obtained in (2) via a suitable choice of the reference function. Nevertheless, it is important to note that in contrast to the aforementioned types of smoothness, anisotropic smoothness is not obtained via a linearization of the cost function around a point and thus it is not straightforward to compare the obtained descent lemmas.
@@ 1.5 Assumptions on the reference function
The reference function $\phi:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is proper, lsc, strongly convex and even with $\phi(0)=0$.
Assumption 1.1 is considered valid throughout the paper. Note that through the duality of strong convexity and Lipschitz smoothness, it implies that $\phi^{*}\in\mathcal{C}^{1}(\mathbb{R}^{n})$. Moreover, from [bauschke2017correction, Proposition 11.7], $\operatorname*{arg\,min}\phi=\{0\}$ and thus $\phi\geq 0$. Throughout the paper we also consider specifically the case where $\phi^{*}\in\mathcal{C}^{2}(\mathbb{R}^{n})$, for which we encode a sufficient condition in the following assumption.
@@
$\operatorname{int}\operatorname{dom}\phi\neq\emptyset$; $\phi\in\mathcal{C}^{2}(\operatorname{int}\operatorname{dom}\phi)$ and for any sequence $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ that converges to some boundary point of $\operatorname{int}\operatorname{dom}\phi$, $\|\nabla\phi(x^{k})\|\to\infty$.
The last condition in Assumption 1.2 is known in the literature as essential smoothness. In light of [rockafellar1977higher, p. 42], under Assumptions 1.1 and 1.2, $\phi^{*}$ is twice continuously differentiable on $\mathbb{R}^{n}$.
@@
When considering a kernel function $h$, in the separable case, it is straightforward that the preconditioner takes the form $\nabla\phi^{*}(x)=({h^{*}}^{\prime}(x_{i}),\dots,{h^{*}}^{\prime}(x_{n}))$. In the isotropic case, the differentiability of $\phi^{*}$ depends on the properties of $h$.
@@
Let $h:\mathbb{R}\to\overline{\mathbb{R}}$ satisfy Assumption 1.1. Then $h^{*}\geq 0$ and $h^{*}$ is an even function, and increasing on $\mathbb{R}_{+}$, while ${h^{*}}^{\prime}(0)=0$. Moreover, $\phi=h\circ\|\cdot\|$ is strongly convex, $\phi^{*}=h^{*}\circ\|\cdot\|$ and
@@
We provide examples of interesting kernel functions in Table 1.
@@ 1.6 Connections with existing methods
As already mentioned, the scheme presented in (2) encompasses various algorithms that are widely used in practice. In this subsection we thus provide some examples that showcase the generalizing properties of our framework.
@@
and by choosing $\lambda=\varepsilon^{-1/2}$ for some $\varepsilon>0$ we retrieve the form of Adagrad [duchi2011adaptive] without memory from [defossez2020simple, Equation (3)] for $\beta_{1}=\beta_{2}=0$.
@@
and by choosing $\lambda=\varepsilon^{-1}$ for $\varepsilon>0$, we retrieve the iterates of Adam [kingma2017adammethodstochasticoptimization, Algorithm 1] where both the exponential decay rates are set to 00, i.e. $\beta_{1}=\beta_{2}=0$.
Another interesting case is that of the clipping gradient method [zhang2019gradient, koloskova2023revisiting].
@@
which is a version of the clipping gradient method.

### v2

# Nonlinearly preconditioned gradient descent under generalized smoothness

@@ Abstract
We analyze nonlinearly preconditioned gradient methods for solving smooth minimization problems. We introduce a generalized smoothness property, based on the notion of abstract convexity, that is broader than Lipschitz smoothness and provide sufficient first- and second-order conditions. Notably, our framework encapsulates algorithms associated with the gradient clipping method and brings out novel insights for the class of $(L_{0},L_{1})$-smooth functions that has received widespread interest recently, thus allowing us to extend beyond already established methods. We investigate the convergence of the proposed method in both the convex and nonconvex setting.
nonconvex optimization,generalized smoothness,first-order methods
@@ 1 Introduction and preliminaries
where $f$ is a continuously differentiable and possibly nonconvex function. While gradient descent is a reliable solver for this type of problems, in many cases it does not fully take advantage of the cost function properties and requires well-tuned or costly stepsize strategies to converge.
In this paper we thus focus on nonlinearly preconditioned gradient methods that are tailored to the properties of the cost function. Given stepsizes $\gamma>0$ and $\lambda>0$ and a starting point $x^{0}\in\mathbb{R}^{n}$, we consider the following iteration:
@@
where $\phi^{*}:\mathbb{R}^{n}\to\mathbb{R}$ is convex and is called the dual reference function. The convex conjugate of $\phi^{*}$, $\phi$ is called the reference function and its properties are crucial to our analysis. This general framework was originally analyzed in (Maddison et al., 2021) under the so-called dual relative smoothness condition for convex and essentially smooth functions. In the general nonconvex and composite nonsmooth setting, it was studied in (Laude & Patrinos, 2025) under a condition called the anisotropic descent property, which was itself first introduced in (Laude et al., 2023) and can be regarded as a globalization of anisotropic prox-regularity (Laude, 2021) for smooth functions.
Our analysis is mainly focused on two types of reference functions that are both generated by a kernel function $h:\mathbb{R}\to\mathbb{R}_{+}\cup\{\infty\}$, thus resulting in two different families of algorithms. The first one is given by the composition with the Euclidean norm, i.e., $\phi=h\circ\|\cdot\|$ and is referred to as the isotropic reference function. In this case, the main iteration (2) takes the form of gradient descent with a scalar stepsize that depends on the iterates. The second is a separable sum obtained via $\phi(x)=\sum_{i=1}^{n}h(x_{i})$, henceforth called the anisotropic reference function. In this case, (2) becomes gradient descent with a coordinate-wise stepsize that depends on the iterates of the algorithm. We remark that although the anisotropic reference function makes the analysis of the method more challenging, it generates more interesting algorithms, akin to the ones that are often used in practice.
@@
[…]
$-|x|-\ln(1-|x|)$ $(-1,1)$ $\frac{y}{1+|y|}$ $\frac{1}{(1+|y|)^{2}}$/✓
$1-\sqrt{1-x^{2}}$ $[-1,1]$ $\frac{y}{\sqrt{1+y^{2}}}$ $(1+y^{2})^{-3/2}$/✓
$x\operatorname{arctanh}(x)-\ln(\cosh(\operatorname{arctanh}(x)))$ $(-1,1)$ $\tanh(y)$ $1-\tanh^{2}(y)$/✓
$\frac{1}{2}x^{2}+\delta_{[-1,1]}(x)$ $[-1,1]$ $\min(1,\max(-1,y))$ $\partial_{C}(\Pi_{[-1,1]})(y)$/
@@ 1.1 Motivation
Unifying framework for clipping algorithms. Gradient clipping and signed gradient methods have garnered attention in recent years due to their efficiency in neural network training and other applications (Bernstein et al., 2018; Gorbunov et al., 2020; Zhang et al., 2020a, b, c; Koloskova et al., 2023; Kunstner et al., 2024). The intuition behind gradient clipping is straightforward, since by clipping one does not allow the potentially very large (stochastic) gradients to hinder the training. Nevertheless, in many cases the clipping threshold and stepsize should be carefully tuned in practice, otherwise leading to suboptimal performance (Koloskova et al., 2023). While algorithms of this type have been analyzed under various smoothness and stochasticity assumptions, there does not seem to exist a simple unifying framework that encapsulates them. Motivated by this gap, we propose a framework that provides further insights into existing methods but also naturally generates new algorithms.
@@
for $\phi=\tfrac{1}{2}\|\cdot\|^{2}$ and $\bar{y}=T_{{L^{-1}},{1}}(\bar{x})=\bar{x}-\tfrac{1}{L}\nabla\phi^{*}(\nabla f(\bar{x}))$. Going beyond the standard Lipschitzian assumptions, it is natural to consider reference functions $\phi$ that generate less restrictive descent inequalities thus allowing us to efficiently tackle more general problems. This is exactly the anisotropic descent property (Laude & Patrinos, 2025, Definition 3.1) and minimizing this upper bound leads to the algorithm described in (2) (for $\lambda$ fixed). Considering (3), it is natural to examine reference functions that are strongly convex and grow faster than a quadratic in order to obtain a less restrictive descent inequality. It turns out that in many interesting cases, the preconditioner in (2) becomes similar to a sigmoid function and the algorithmic step takes the form of popular algorithms. We present examples of kernel functions and the corresponding preconditioners in Table 1.
The abstraction discussed in the previous paragraph is in fact tightly connected to the notion of $\Phi$-convexity (also known as $c$-concavity in the optimal transport literature), which states that a function is $\Phi$-convex if it can be written as the pointwise supremum over a family of nonlinear functions. Similarly to the fact that every proper, lsc and convex function can be expressed as a pointwise supremum of its affine minorizers (Rockafellar & Wets, 1998, Theorem 8.13), anisotropic smoothness then requires that $-f$ is a pointwise supremum of nonlinear minorizers. This fact is once again in parallel to classical $L$-smoothness, which requires that $-f$ is the pointwise supremum over concave quadratics, and leads to an envelope representation of $f$ that is useful in studying the corresponding calculus. In that regard, anisotropic smoothness is a straightforward extension of Lipschitz smoothness. A visualization of the concept of $\Phi$-convexity is shown in Figure 1.
@@ 1.2 Our contribution
  • We describe a common nonlinear gradient preconditioning scheme for the main iterates, i.e., without momentum nor exponential moving average mechanisms, of popular algorithms including gradient clipping, Adam, Adagrad and recently introduced methods for $(L_{0},L_{1})$-smoothness. These preconditioners are gradients of smooth convex functions and have sigmoid shape, reminiscent of common activation functions in neural networks.
  • We introduce $(L,\bar{L})$-anisotropic smoothness, which extends (Laude & Patrinos, 2025) allowing for two constants and a reference function $\phi$ of possibly non-full domain and prove that it is less restrictive than $L$-smoothness. Through a novel technique we study necessary and sufficient first- and second-order conditions for $(L,\bar{L})$-anisotropic smoothness, in the process obtaining novel characterizations for the forward operator $T_{{\gamma},{\lambda}}$, which in turn leads to new insights into the class of $(L_{0},L_{1})$-smooth functions.
  • We analyze the convergence of (2) in the nonconvex setting, obtaining new results for our stationarity measure. In the convex setting, we prove the sublinear rate of the method for a large family of isotropic reference functions, utilizing only a simple dual characterization. In the more challenging case where $\phi$ is anisotropic, we present an unconventional proof that is based on the envelope representation of anisotropic smoothness. We are thus able to obtain standard $O(1/K)$ convergence rates for large classes of functions.
@@ 1.3 Related work
Dual space preconditioning and anisotropic smoothness. The scheme described in (2) was originally introduced in (Maddison et al., 2021) in the convex setting, where it was analyzed under a condition called dual relative smoothness for which sufficient second-order conditions were provided. In (Laude & Patrinos, 2025) the anisotropic smoothness condition was studied, which was shown to naturally lead to the convergence of the method in the nonconvex and proximal case. Moreover, in (Léger & Aubin-Frankowski, 2023) the scheme was also analyzed under the general framework of $\Phi$-convexity and a sufficient second-order condition for anisotropic smoothness was provided. Nevertheless, this requires that $\phi\in\mathcal{C}^{4}(\mathbb{R}^{n})$ satisfies the non-negative cross curvature condition from optimal transport (NNCC) (see (Figalli et al., 2011, Assumption (B3)) and (Léger & Aubin-Frankowski, 2023, Definition 2.8)), which is a strong assumption that does not hold for many interesting reference functions. An accelerated version of the method for Lipschitz smooth problems was introduced and studied in (Kim et al., 2023). Recently, the method was also extended to measure spaces in (Bonet et al., 2024). Furthermore, a relaxed proximal point algorithm with nonlinear preconditioning akin to (2) for solving monotone inclusion problems was studied in (Laude & Patrinos, 2023).
@@
Generalized smoothness. Our work is also connected to other notions of generalized smoothness, i.e., descent inequalities beyond the standard Lipschitzian assumptions. To begin with, Bregman (relative) smoothness is a popular extension of Lipschitz smoothness (Bauschke et al., 2017a; Lu et al., 2018; Bolte et al., 2018; Ahookhosh et al., 2021) that can encapsulate a wide variety of functions such as those whose Hessians exhibit a certain polynomial growth (Lu et al., 2018). Other notions of generalized smoothness include Hölder smoothness (Bredies, 2008; Nesterov, 2015) and also higher-order smoothness where higher-order derivatives of $f$ are Lipschitz continuous (Nesterov & Polyak, 2006; Doikov & Nesterov, 2020).
@@
Recently, a new concept of smoothness has been introduced in order to capture the cases where the norm of the Hessian is upper bounded by some function of the norm of the gradient (Zhang et al., 2020b; Chen et al., 2020; Li et al., 2024). This condition, which in its most popular form is called $(L_{0},L_{1})$-smoothness (Zhang et al., 2020b, Definition 1), has received widespread attention. Various existing methods have been analyzed under this new smoothness condition (Wang et al., 2023; Faw et al., 2023; Koloskova et al., 2023), while also new ones have been proposed (Gorbunov et al., 2024; Vankov et al., 2024). We remark that a number of these algorithms can actually be obtained in (2) via a suitable choice of the reference function. Nevertheless, it is important to note that in contrast to the aforementioned types of smoothness, anisotropic smoothness is not obtained via a linearization of the cost function around a point and thus it is not straightforward to compare the obtained descent lemmas.
@@ 1.4 Notation
Let $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a locally Lipschitz function, we denote the (Clarke) generalized Jacobian as $\partial_{C}F(x)=\operatorname{con}\{\lim_{x_{i}\to x}\nabla F(x_{i}):x_{i}\notin\Omega_{F}\}$, where $\Omega_{F}$ is the set of points where $F$ fails to be differentiable. $\Pi_{C}$ denotes the projection on a set $C$. For an $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ we say that it is $(L_{0},L_{1})$-smooth for some $L_{0},L_{1}>0$ if it holds that $\|\nabla^{2}f(x)\|\leq L_{0}+L_{1}\|\nabla f(x)\|$ for all $x\in\mathbb{R}^{n}$. Otherwise we adopt the notation from (Rockafellar & Wets, 1998).
@@ 1.5 Assumptions on the reference function
The reference function $\phi:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is proper, lsc, strongly convex and even, i.e., $\phi(x)=\phi(-x)$, with $\phi(0)=0$.
Assumption 1.1 is considered valid throughout the paper. Note that through the duality of strong convexity and Lipschitz smoothness, it implies that $\phi^{*}\in\mathcal{C}^{1}(\mathbb{R}^{n})$. Moreover, from (Bauschke et al., 2017b, Proposition 11.7), $\{0\}=\operatorname*{arg\,min}\phi$ and thus $\phi\geq 0$. Throughout the paper we also consider specifically the case where $\phi^{*}\in\mathcal{C}^{2}(\mathbb{R}^{n})$, for which we encode a sufficient condition in the following assumption, in light of (Rockafellar, 1977, p. 42).
@@
$\phi\in\mathcal{C}^{2}(\operatorname{int}\operatorname{dom}\phi)$ is essentially smooth.
It is important to note that through (Rockafellar, 1977, p. 42), under Assumptions 1.1 and 1.2, the Hessian matrix of $\phi^{*}$ is positive-definite everywhere.
@@
When considering a kernel function $h$, in the anisotropic case, it is straightforward that $\phi$ inherits the properties of $h$ and the preconditioner takes the form $\nabla\phi^{*}(x)=({h^{*}}^{\prime}(x_{1}),\dots,{h^{*}}^{\prime}(x_{n}))$. In the isotropic case, the differentiability of $\phi^{*}$ depends on the properties of $h$, as we show next.
@@
Let $h:\mathbb{R}\to\overline{\mathbb{R}}$ satisfy Assumption 1.1. Then $h^{*}\geq 0$ is an even function and increasing on $\mathbb{R}_{+}$, while ${h^{*}}^{\prime}(0)=0$. Moreover, $\phi=h\circ\|\cdot\|$ is strongly convex, $\phi^{*}=h^{*}\circ\|\cdot\|$ and
@@
We provide examples of interesting kernel functions, along with the assumptions that they satisfy, in Table 1.
@@ 1.6 Connections with existing methods
As already mentioned, the scheme presented in (2) encompasses the basic iterations of various algorithms that are widely used in practice. In this subsection we thus provide some examples that showcase the generalizing properties of our framework.
@@
and by choosing $\lambda=\varepsilon^{-1/2}$ for some $\varepsilon>0$ we retrieve the form of Adagrad (Duchi et al., 2011) without memory from (Défossez et al., 2022, Equation (3)) with $\beta_{1}=\beta_{2}=0$.
@@
and by choosing $\lambda=\varepsilon^{-1}$ for $\varepsilon>0$, we retrieve the iterates of Adam (Kingma & Ba, 2014, Algorithm 1) where both the exponential decay rates are set to 00, i.e. $\beta_{1}=\beta_{2}=0$.
@@
$x^{k+1}=x^{k}-\min(\gamma/\|\nabla f(x^{k})\|,\gamma\lambda)\nabla f(x^{k}).$
Note that the gradient clipping method as presented in (Zhang et al., 2020b, Equation (5)) is given by $x^{k+1}=x^{k}-\min\{\eta_{c},\tilde{\gamma}\eta_{c}/\|\nabla f(x^{k})\|\}\nabla f(x^{k})$. Therefore, by choosing $\gamma=\tilde{\gamma}\eta_{c}$ and $\lambda=1/\tilde{\gamma}$ we can see that (2) encompasses the gradient clipping method.